### v1

# IndRegBias: A Dataset for Studying Indian Regional Biases in English and Code-Mixed Social Media Comments

Abstract
Warning: This paper consists of examples representing regional biases in Indian regions that might be offensive towards a particular region.
While social biases corresponding to gender, race, socio-economic conditions, etc., have been extensively studied in the major applications of Natural Language Processing (NLP), biases corresponding to regions have garnered less attention. This is mainly because of (i) difficulty in the extraction of regional bias datasets, (ii) disagreements in annotation due to inherent human biases, and (iii) regional biases being studied in combination with other types of social biases and often being underrepresented. This paper focuses on creating a dataset IndRegBias, consisting of regional biases in an Indian context reflected in users’ comments on popular social media platforms, namely Reddit and YouTube. We carefully selected 25000 comments appearing on various threads in Reddit and videos on YouTube discussing trending topics on regional issues in India. Furthermore, we propose a multilevel annotation strategy to annotate the comments describing the severity of regional biased statements. To detect the presence of regional bias and its severity in IndRegBias, we evaluate open-source Large Language Models (LLM) and Indic Language Models (ILM) by using zero-shot, few-shot, and fine-tuning strategies. We observe that zero-shot and few-shot approaches show lower accuracy in detecting regional biases and severity in the majority of the LLMs and ILMs. However, the fine-tuning approach significantly enhances the performance of the LLM in detecting Indian regional bias along with its severity.
IndRegBias: A Dataset for Studying Indian Regional Biases in English and Code-Mixed Social Media Comments
Debasmita Panda1, Akash Anil1, Neelesh Kumar Shukla2 1Department of Data Science and Engineering, Indian Institute of Science Education and Research, Bhopal 2Oracle Industries AI, Oracle Corporation {debasmitap21, anila}@iiserb.ac.in, neelesh.kumar.shukla@oracle.com
1 Introduction
Natural Language Processing (NLP) has shown an impressive footprint in various applications such as Question-Answering Devlin et al. (2019), Content Summarization See et al. (2017), Search Engines Mitra et al. (2018), etc. NLP-based models use a large number of real-world texts in modelling the underlying task Bender et al. (2021). Consequently, NLP models are subject to various biases related to social Sap et al. (2019), geographic Faisal and Anastasopoulos (2023), linguistic Fleisig et al. (2024), and cultural Hershcovich et al. (2022) factors, which may result in biased predictions Zhao et al. (2019). Thus, in the recent past, there has been a surge in studying various biases Bolukbasi et al. (2016); Dixon et al. (2018); Sun et al. (2019) for popular NLP-based models such as Large Language Models (LLMs).
The majority of the biases, such as linguistics and culture, appear due to either over-representation or under-representation of data Hovy and Spruit (2016). However, social bias can be evident due to common societal stereotypes regarding race Dixon et al. (2018), gender Bolukbasi et al. (2016), economics, and region Jardim and others (2022). Although social biases based on gender, race, and economics have been studied extensively Sun et al. (2019), only a few studies consider social biases corresponding to regions Bhatt et al. (2022); Jha and Mamidi (2023). However, regional stereotypes are prevalent in many diverse countries, including India, as recorded from multiple news media111(1)Northeast citizens faced racial discrimination amid COVID-19 outbreak ; (2) The Great Indian Bias ; (3) Bengaluru man demands North Indian be kept out of the city. Thus, this work focuses on revisiting social bias due to regional stereotypes, referred to as Regional bias (RB) in the subsequent part of the paper.
Language(s) Comment Translation
English No offence, but Kannadigas really need to chill about language. It’s not that deep. No offence, but Kannadigas really need to chill about language. It’s not that deep.
Hindi Arre Bihari log ka logic sunke hi dimaag hil jaata hai ’Humko sab aata hai’ waale IAS ban jaate hain kaise bhai? Just hearing Bihari people’s logic shakes my brain. How do these ’we know everything’ types become IAS officers?
Marathi, English, Hindi Marathi manoos ka ek hi kaam – complain about outsiders and vote for MNS. Kuch kaam bhi kar lo bhai. The only thing a Marathi person does is complain about outsiders and vote for MNS. Do some actual work too, bro.
Table 1: Examples of Indian regional bias expressed through different languages and sarcasm.
Figure 1: Construction of IndRegBias and detecting regional bias with severity using Large Language Models and Indic Language Models. (In the annotation phase, if a comment is regional bias, then the second and third labels are determined. Otherwise, the comment gets a single label of 0. For example, after the data annotation phase, a text/comment is assigned labels as (1, 2, L) where 1 is for stating the comment is regional bias, 2 states the bias is Moderate, and L states the region being targeted.
Regional bias broadly refers to cognitive prejudices or stereotypical assumptions directed toward individuals based on their regional identity. Table 1 presents comments from popular social media, namely Reddit [^2] https://www.reddit.com/ and YouTube [^3] https://www.youtube.com/ by users towards Indian states or regions. It is evident that these texts carry stereotypes about natives of different regions and thus promote regional bias. Previous works Bhatt et al. (2022); Jha and Mamidi (2023) focused on benchmarking RBs written in English. However, it is evident from Table 1 that detecting RB in NLP models requires semantic understanding of the complete text, often written in transliterated and code-mixing of multiple languages. Thus, this paper focuses on creating a novel text dataset capturing regional biases in India, incorporating 36 regional boundaries [^4] Defined by the States Reorganization Act and subsequent amendments. referred to as state, capital territory, and union territory possessing diverse regional characteristics Office of the Registrar General & Census Commissioner, India (2011).
To construct the dataset of regional bias in the Indian context IndRegBias, this paper exploits comments from Reddit and YouTube. We chose social media comments because, due to anonymity and less regularization on social media, users tend to express their real feelings as comments corresponding to other regions Suler (2004). Figure 1 presents the data generation and evaluation pipeline. In the data generation phase, we selected some of the recent topics on regional tensions in India and scraped 25000 user comments expressing their sentiments and beliefs on these topics over Reddit and YouTube. Two groups, each of three university students having diverse backgrounds and knowledge of at least three Indian languages, meticulously annotated the comments using a multilevel annotation framework resulting in three labels for a comment, i.e., given a comment $\mathcal{C}$, the first label states whether $\mathcal{C}$ is a regional bias or not ?, the second label gives the severity of $\mathcal{C}$ (Mild: 1, Moderate: 2, and Severe: 3) given $\mathcal{C}$ is a regional bias, and the third label states which region (state or union/capital territory) is being targeted given $\mathcal{C}$ is a regional bias.
Recently, LLMs such as GPT Achiam et al. (2023), LLaMa Touvron et al. (2023), and DeepSeek Liu et al. (2024) have emerged as efficient frameworks for carrying out multiple applications, e.g., question-answering and recommender systems. These frameworks are used by millions of users daily for their day-to-day queries. Most of these LLMs have been trained considering explicit knowledge of harmful content such as hate speech, slang, and other derogatory words, which is evident from their curated and restricted results. However, detecting regional biases in Indian texts remains a significant challenge due to the nation’s diverse cultures, languages, and beliefs  Bhatt et al. (2022). While large language models (LLMs) have demonstrated strong performance across many tasks, they are not inherently attuned to the subtle and context-dependent aspects of Indian regional biases. This limitation arises because LLMs are trained on broad datasets where nuanced or underrepresented regional characteristics may not be adequately captured. To address this, as depicted in the LLM evaluation pipeline in Figure  1, we systematically evaluated LLMs using zero-shot, few-shot, and fine-tuning experimental setups. These settings enable us to assess how well LLMs can generalize to the task of regional bias detection (binary classification) and severity ranking (multiclass classification), and to what extent their performance improves with domain-specific examples or further adaptation.
The rest of the paper is organized as follows: Section 2 presents the related studies on bias and NLP. Section 3 describes the proposed regional biased dataset in the Indian context, followed by Section 4 presenting the experimental setups for LLM evaluation. Section 5 presents the results and discussion. Finally, Section 6 concludes the paper.
2 Related Studies
Bias in NLP and Regional Bias
Social biases in NLP models have primarily been studied in the context of gender Bolukbasi et al. (2016); Sun et al. (2019), race Dixon et al. (2018), and religion Abid et al. (2021). These works have demonstrated that a majority of the language models encode the inherent stereotypes and may result in biased prediction Bender et al. (2021). While regional identity is an important factor under the social system, less attention has been paid to studying and curating resources corresponding to regional biases Hovy and Prabhumoye (2021). Recently, Adebayo and others (2024) discusses the regional biases in monolingual English language models and contributes an investigation framework specifically tailored for low-resource regions, revealing that there are multiple geographic variations in the word embeddings by BERT Devlin et al. (2019) while they were assumed to be similar. Further, Borah and others (2025) identifies topical differences in gender bias across different regions and proposes a region-aware bottom-up approach for bias assessment. To measure the hierarchical regional bias, Li and others (2022) proposes a metric HERB that utilizes the information from the sub-region clusters to quantify the bias in pre-trained language models.
While these studies consider regional aspects, they are limited to analytics supported by other types of biases.
Regional Bias in Multilingual Indian Regions
Regional biases in Indian regions are commonly prevalent and are often seen in terms of dialect variations Kumar and others (2021), transliteration Roark and others (2012), and code-mixed uses Bali and others (2014). In a recent work, Bhatt et al. (2022) released a fairness evaluation corpus covering stereotypes pertaining to region and religion axes relevant to the Indian context in the English language. Further, Jha and Mamidi (2023) introduces a novel benchmark SeeGULL in English containing stereotypes about identity groups spanning 178 countries across 8 different geopolitical regions across 6 continents, as well as state-level identities within the US and India using generative models. While these resources are related to Indian regions, they have been curated in English, which might lose the inherent bias characteristics of the local regions within India.
Although not specific to regional bias, Sahoo and others (2024) introduces IndiBias, a benchmark dataset designed particularly for evaluating social biases in the Indian context and capturing different variants of social biases in the Hindi and English languages.
These datasets bridge the gap in social biases with Indian contexts, but as discussed, regional biases in real-world communication are more nuanced and often found transliterated or code-mixed. Thus, our paper focuses on introducing IndRegBias, consisting of real-world discussions (using comments from social media) over regional issues, which can bridge the gap for regional bias datasets in the Indian context.
3 Creation Of IndRegBias
The creation of IndRegBias is motivated by the scarcity of Indian benchmark datasets, particularly for regional biases. While the other benchmarks created in Sahoo and others (2024); Jha et al. (2023) capture Indian contexts, they are based on translation primarily from the English language. However, with the freedom of speech on social media, a majority of users post their stereotypes anonymously Suler (2004). We argue that NLP-based models should be robustly trained over such kinds of texts to promote regional inclusivity.
3.1 Data Collection
To prepare regional bias datasets over social media, we consider two popular online discussion platforms, namely, Reddit and YouTube. We selected subreddit pages focusing on discussions about the stereotypes and biases targeting particular regions and states in India. Similarly, we selected YouTube channels that specifically discuss individual states, the biases related to them, the prevalent problems in the state, the issues associated with the identity of the people, and interviews with residents.
We extracted the comments using praw (Python Reddit API Wrapper) Boe (2024) for Reddit and googleapiclient Google Developers (2024) for YouTube. The comments collected from these platforms contain numerous spam comments that are irrelevant to our work. Thus, after cleaning and preprocessing, a final dataset of 25,000 comments was curated for annotation. Further, as the comments were collected from videos or subreddit pages belonging to different Indian regions, where the languages are mixed, such as English, Hinglish (a mix of Hindi and English), a mix of Bengali and English, a mix of Malayalam and English, or Marathi and English, etc., it results in a transliterated and code-mixed dataset naturally representing the Indian diversity.
3.2 Data Annotation
The nuances in social media comments, particularly in the Indian context, make the annotation task very challenging. Thus, we propose a multilevel annotation policy to ensure high-quality labeling along with maintaining consistency.
3.2.1 Multilevel Annotation Policy
The goal of the annotation task was to categorize social media comments into regional biases or non-regional biased categories. The annotation process was divided into three sub-tasks:
Task 1: Bias Identification
Each comment was classified into one of two categories:
  • Regional Bias (RB: 1): The comment propagates a stereotype (positive or negative) regarding a state, region, caste, politics, or culture associated with a specific Indian region.
  • Non-Regional Bias (NRB: 0): The comment does not exhibit stereotypes or is unrelated to the objective.
Task 2: Severity Scoring The comments classified as RB in Task 1 were assigned a severity score based on the definitions in Table 2. Positive stereotypes were explicitly mapped to the “Mild” category.
Score Level Definition
0 None No regional bias present.
1 Mild Slight bias, satire, or positive stereotypes (e.g., “Himachalis are very simple”).
2 Moderate Clear bias reinforcing negative stereotypes with intent to mock.
3 Severe Strong bias promoting discrimination, racism, or dehumanization.
Table 2: Severity scoring criteria for regional bias.
Task 3: Target Region Assignment Annotators identified the target(s) of the comment from a predefined list of regions (e.g., North-India, NorthEast-India, Central India, etc.) and 30+ specific States/Union Territories (e.g., Punjab, Tamil Nadu, etc.).
[…]
$\kappa=\frac{p_{o}-p_{e}}{1-p_{e}}$ (1)
where $p_{o}$ represents the relative observed agreement among raters, and $p_{e}$ represents the hypothetical probability of chance agreement. Upon calculation, the Cohen’s kappa value for binary classification (RB vs NRB) came to be 0.91, and 0.83 for multi-class classification (severity and region assignment combined), indicating strong agreement. Further disagreements were resolved through comprehensive discussions among the annotators, including one neutral annotator. Table 3 demonstrates a sample of annotated IndRegBias.
Comment Is RB? Severity state/region
“Biharis have a weird accent.” 1 1 Bihar
“Jharkhand is just a jungle.” 1 2 Jharkhand
“North Indians are just casteist and illiterate” 1 3 North-India
“Himachalis are so sweet.” 1 1 Himachal Pradesh
“Proud to be an Assamese.” 0 0 NA
Table 3: A sample of IndRegBias with fine-grained annotation using the multi-level annotation policy
3.3 Data Analysis
After annotation of 25000 comments, IndRegBias consists of 13,015 (52.1%) regional biases, while 11,985 (47.9%) are non-regional biases. Under the severity label for regional biased comments approximately 29% are Mild, 51% are Moderate, and 20% are Severe.
3.3.1 Data Distribution based on Region
Figure 2 presents the distribution of regional biases IndRegBias categorized into 6 Indian regions (with overlapping possible if a comment targets more than one region), namely North, South, East, West, NorthEast, and Central (based on collections of some of the states that are closer geographically). We observe that approximately 40% of biased comments target South Indians followed by North India with 26%. Further, East and West Indian regions are targeted by approximately 16% and 10% of comments, respectively. We record a lower number of comments, approximately 5% and 2% of comments targeting NorthEast and Central region respectively.
3.3.2 Data Distribution based on States/Territories
Figure 3 shows the distribution of regional biases categorized in 29 Indian states. We observe that the distribution loosely describes the long-tailed distribution where few of the states, such as Kerala, Goa, West Bengal, Karnataka, Bihar, and Gujarat, are targeted approximately 60% of the time.
05005001,0001{,}0001,5001{,}5002,0002{,}0002,5002{,}5003,0003{,}0003,5003{,}5004,0004{,}0004,5004{,}5005,0005{,}0005,5005{,}5006,0006{,}0006,5006{,}5007,0007{,} [^0] Central-IndiaNorthEast-IndiaWest-IndiaEast-IndiaNorth-IndiaSouth-India5,8205{,}8203,8443{,}8442,3072{,}3071,5081{,}508827827339339Number of Comments
Figure 2: Distribution of data across different regions in India.
02002004004006006008008001,0001{,}0001,2001{,}2001,4001{,}4001,6001{,}6001,8001{,}8002,0002{,}0002,2002{,}2002,4002{,}400TripuraMeghalayaManipurSikkimMizoramArunachal Pr.AssamUttarakhandAndhra Pr.NagalandTelanganaHimachal Pr.OdishaChhattisgarhMadhya Pr.JharkhandHaryanaRajasthanJ&KPunjabMaharashtraUttar PradeshTamil NaduGujaratBiharKarnatakaWest BengalGoaKerala12121616272727273030323249496565727298981061061071071111111631631731732102102172172182182182184124124384385995996146148328328718718878871,0911{,}0911,0951{,}0951,8571{,}857Number of Comments
Figure 3: Distribution of regional biased comments across Indian states.
3.4 Qualitative Analysis of Regional Biased Comments
Theme States/Regions Stereotypes
Crime & Corruption Delhi, Uttar Pradesh, Tripura, J&K, Goa Frequently labelled with stereotypes related to high crime rates, corruption, drug trafficking, and organised crime.
Social Intolerance Delhi, Mizoram, Karnataka, Haryana, Uttar Pradesh, J&K, Manipur Stereotypes concerning racism, arrogance, misogyny, casteism, and anti-national sentiments.
Socio-Economic Failure Bihar, Uttar Pradesh, Jharkhand Branded as “failed” or “backward,” focusing on poverty, unemployment, illiteracy, and caste conflict (framed as a “national burden”).
Table 4: Categorization of severe regional bias comments
010102020303040405050606070708080GoaJ&KTripuraManipurHaryanaBiharJharkhandKarnatakaNorth-IndiaDelhiUttar PradeshMizoram282831313333303031313030373725252222333330302323Number of Comments
Figure 4: Distribution of severe bias comments across different states.
We perform qualitative analysis of IndRegBias, which resulted in distinct patterns for the severity of the biases present in the comments, having aggression and using derogatory language frequently to describe their ideas over the states or the people of the state. We categorized the states having severe regional biases into three main themes, namely Crime and Corruption, Social Intolerance, and Socio-Economic Failures, as summarized in Table 4. Figure 4 shows the distribution of severe comments under the three themes. We observed that these three themes share approximately a similar number of comments when distributed qualitatively. For example, Delhi received 33 comments under Crime and Corruption and Social Intolerance, and Jharkhand and Bihar received 37 and 30, respectively.
4 Experimental Setups and Analysis
The experimental analysis begins with evaluating the models in zero-shot and few-shot settings, followed by fine-tuning to improve their identification of regional biases. For zero-shot Kojima et al. (2022) and few-shot Brown et al. (2020), we have used the chain-of-thought prompting Wei et al. (2022) technique described in Section 9.1 to classify comments into binary or multi-class labels.In the few-shot setting, we have conducted experiments across different support settings, of which the support was selected randomly, and inferencing was conducted on the rest of the comments. The rationale was provided in Section 9.2. For the fine-tuning experiment, we have employed Parameter-Efficient Fine-Tuning (PEFT) Xu et al. (2023) using Low-Rank Adaptation (LoRA) Hu et al. (2021). This approach freezes the pre-trained weights and injects trainable adapters, significantly reducing computational overhead. We have used a random dataset split of 70% Train, 10% Validation, and 20% Test. To ensure robust results and mitigate overfitting, we have implemented 5-Fold Stratified Cross-Validation. We have applied two fine-tuning techniques: (i) Instruction-Based Supervised Fine Tuning (SFT) for binary classification for detecting a comment as regional bias or not, and Classification-Based SFT for multi-class classification, where we replace the generative layer with a classification head to predict severity levels (1–3) directly. A WeightedRandomSampler is employed to address class imbalance. All hyperparameters for the fine-tuning experiments are listed in Appendix Section 9.3.
5 Results and Discussion
We conducted systematic experiments to evaluate LLMs and ILMs for the detection of regional biases in IndRegBias using zero-shot, few-shot, and fine-tuning approaches. [^5] All the prompts are available in the Appendix. The dataset and codes are available on GitHub. At first, detecting regional bias is achieved by employing the language models as binary classifiers. Additionally, to assess severity, we employ a multi-class classification method wherein the language models predict a severity class (i.e., Mild, Moderate, and Severe). We utilize conventional machine learning measures to evaluate the performance of different models.
5.1 Regional Bias (RB) vs Non-Regional Bias (NRB): Binary Classification
Table 5 presents the performance of 11 open-source large language models (8 LLMs with variants + 3 ILMs) for identifying comments as RB or NRB using zero-shot inference [^6] A small snapshot of IndRegBias consisting of 5000 comments was evaluated using proprietary model GPT-4o and LLama 4 Scout. We observe that open-source models such as Qwen were achieving similar performance. Table 9.4 gives the results of this setting in the Appendix. It is evident that Qwen3 (8B and 32 B) consistently outperformed all other LLMs. This might be due to extensive multilingual pre-training of Qwen with 119 languages, including Indian languages Team (2025), capturing the regional semantics. Moreover, Krutrim-2 outperforms all other ILMs and maintains competitiveness in comparison to Qwen3. This may be attributed to Krutrim-2’s utilization of Mistral-Nemo and its targeted pre-training in 13 Indian languages Kallappa et al. (2025).
Inferencing the above language models in the few-shot setting is costly in terms of GPU requirements and time. Thus, we select the Qwen3-8B model from Table 5, as it is competitive to Qwen3-32B but requires less memory. We first do a preliminary experiment on a smaller snapshot for selecting supports in a few-shot evaluation [^7] Refer to Section 9.8 in Appendix. As shown in Table 6, providing 50 regional bias examples (Exp-1) yielded the most improvement, boosting the F1-score to 0.82 over the zero-shot scores. The balanced support set (Exp-2) attained the highest Precision (0.85) but experienced a reduction in F1-score (0.75), signifying a notable trade-off in recall. Exp-3 improved both Precision (0.80) and F1-score (0.81).
As discussed in Section 4, we select the splits 70%, 10%, and 20% for IndRegBias and evaluate the performance of fine-tuning Qwen3-8B and Qwen3-32B in a five-fold stratified cross-validation setting. Table 7 presents the average of fine-tuning results compared with the zero-shot performance of the same testing examples. It is evident that fine-tuned Qwen3-8B and Qwen3-32B models consistently achieved better performance than the zero-shot setting. This justifies the requirement of a focused dataset to train the existing LLMs for capturing the regional biases.
Table 5: Zero-Shot Binary Classification for Regional Bias Detection in IndRegBias. [Acc: Accuracy, P: Precision, F1: F1-Score]
Model Acc P F1
Qwen3-8B 0.74 0.71 0.78
Qwen3-32B 0.74 0.79 0.72
Krutrim-2 0.73 0.68 0.78
Mistral-7B-v0.3 0.70 0.65 0.76
Gemini-2.5-Pro 0.69 0.64 0.76
Airavata 0.58 0.60 0.58
Sarvam-M 0.57 0.63 0.49
Llama-3.2-3B 0.55 0.57 0.57
DeepSeek-R1 0.54 0.53 0.69
Gemma-1.1-7B-IT 0.52 0.52 0.69
Mistral-Nemo 0.53 0.53 0.69
Phi-4-Mini 0.48 0.60 0.00
Table 6: Few-Shot (FS) vs Zero-Shot (ZS) Performance by Qwen3-8B. [R: Regional Bias, N: Non-Regional Bias]
Exp-1 Exp-2 Exp-3
(50 R) (25R/25N) (30R/20N)
Metric ZS FS ZS FS ZS FS
Precision 0.70 0.74 0.71 0.85 0.71 0.80
F1-Score 0.77 0.82 0.77 0.75 0.78 0.81
Table 7: Comparison of Fine-Tuning (FT) vs. Zero-Shot (ZS) for Qwen3 Models.
Qwen3-8B Qwen3-32B
Metric ZS FT ZS FT
Precision 0.692 0.902 0.790 0.900
F1-Score 0.772 0.904 0.724 0.902
5.2 Detecting Severity of Regional Bias: Multi-Class Classification
We selected the four best-performing models from Table 5 for zero-shot multi-class severity classification of the regional bias comments. Table 8 presents the result of severity prediction using the zero-shot approach. It is evident that Mistral-7B-v0.3 achieved the highest Precision for the ‘Mild’ and ‘Moderate’ classes (0.64) and the best F1-score for the ‘Severe’ class (0.45). Qwen3-8B excelled for the ‘Severe’ class with Precision (0.44) and for the ‘Moderate’ class with F1-score (0.66). Although Mistral-Nemo-Base performed comparably to Qwen3-8B, Qwen3-32B lagged in both the ‘Mild’ and ‘Severe’ categories.
Table 8: Zero-Shot Multi-class severity detection for Regional Biases. [P: Precision, F1: F1-Score]
Mild Mod. Severe
Model P F1 P F1 P F1
Qwen3-8B 0.57 0.57 0.59 0.66 0.44 0.22
Qwen3-32B 0.51 0.58 0.62 0.57 0.39 0.36
Mistral-7B 0.64 0.62 0.64 0.54 0.36 0.45
Mistral-Nemo 0.52 0.58 0.61 0.59 0.41 0.35
Krutrim-2 0.48 0.59 0.61 0.55 0.41 0.32
For few-shot multi-class classification, we evaluate the best-performing model (Mistral-7B-v0.3) using three support configurations. As shown in Table 9, performance relative to the zero-shot baseline is mixed. The larger balanced set (Exp B) improved stability, significantly boosting the ‘Moderate’ F1-score (0.54 to 0.64) and ‘Severe’ Precision (0.36 to 0.41), though at the cost of ‘Severe’ recall.
Table 9: Few-Shot Multi-class Classification Results compared to Zero-Shot baselines.[P: Precision, F1: Score]
Zero-Shot Few-Shot
Class P F1 P F1
Exp A: 9 Examples (3 per class)
Mild (1) 0.64 0.62 0.58 0.62
Moderate (2) 0.64 0.54 0.64 0.58
Severe (3) 0.36 0.45 0.40 0.44
Exp B: 30 Examples (10 per class)
Mild (1) 0.64 0.62 0.66 0.61
Moderate (2) 0.64 0.54 0.65 0.64
Severe (3) 0.36 0.45 0.41 0.41
Exp C: 30 Examples (Imbalanced)
Mild (1) 0.64 0.62 0.62 0.63
Moderate (2) 0.64 0.54 0.62 0.67
Severe (3) 0.35 0.45 0.48 0.31
For fine-tuning in the multi-class severity detection setting, we fine-tuned Mistral-7B-v0.3 using 5-fold stratified cross-validation on regional bias comments only. The results presented in Table 10 demonstrate significant improvements over zero-shot across all classes. Notably, the ’Severe’ category has considerable gains, validating the effectiveness of fine-tuning on our dataset.
Table 10: Zero-Shot vs. Fine-Tuning Performance (Mistral-7B-v0.3).
Zero-Shot Fine-Tuning
Severity Level P F1 P F1
Mild (1) 0.79 0.62 0.76 0.68
Moderate (2) 0.63 0.61 0.66 0.72
Severe (3) 0.36 0.45 0.59 0.53
6 Conclusion
In this paper, we introduced IndRegBias, a novel dataset comprising 25,000 user comments from social media discussions on Indian regional issues. All comments were annotated by human experts following a multi-level labeling policy to assess both the presence of regional bias and its severity, resulting in a balanced dataset of regional and non-regional bias examples. We conducted a systematic evaluation of large language models (LLMs) and Indic language models under zero-shot, few-shot, and fine-tuning experimental settings. Our results indicate that off-the-shelf even the well advanced models exhibit limited ability to detect and rank nuanced regional biases, particularly in the zero-shot setting; however, performance improves substantially with fine-tuning on domain-specific data. This work highlights a critical gap in existing LLM knowledge and provides a valuable resource for developing more culturally informed and region-aware language models.
7 Limitations
While IndRegBias serves as a robust benchmark for Indian regional bias, we acknowledge the limitations of our data collection and evaluation methodology.
Geographic and Demographic Skew:
As detailed in Section3.3, the dataset shows a geographic imbalance. Regions such as South India ($40\%$) and North India ($26\%$) are over-represented compared to Northeast India ($5\%$) and Central India ($2\%$). This imbalance is major across the states as well, which needs to be addressed.
Subjectivity in Annotation:
Despite a rigorous multi-level annotation policy and high inter-annotator agreement ($\kappa=0.83$), the perception of “Severity” remains subjective. The distinction between “Mild” and “Moderate” can be heavily influenced by the annotator’s cultural background and familiarity with specific regional context.
Model Safety Refusals:
A significant limitation in evaluating proprietary or highly aligned LLMs (e.g., Gemini-2.5-Pro, Llama-3.2, Sarvam-M) is their tendency to trigger safety refusals. As stated in Appendix 9.7, these models frequently decline to classify text that includes slurs or derogatory stereotypes, resulting in a neutral output or a refusal string. This behavior lowers their quantitative performance (Recall) on our benchmark, not necessarily due to a lack of understanding, but due to rigid safety alignment that treats analysis of hate speech as generation of hate speech.
8 Ethics Consideration
Data Collection and Privacy:
The dataset IndRegBias consists of 25,000 comments collected from social media platforms (Reddit and YouTube) as detailed in 3. We have utilized the official APIs (PRAW and Google API Client) to ensure compliance with the platforms’ terms of crawling. To protect the user privacy, all personally identifiable information, such as username, userId, and profile URLs, has been removed in the preprocessing stage. The data collection was purely observational, focusing exclusively on the public interactions related to regional biases, without influencing discussions or engaging with users.
Human Annotation and Quality Control:
Human annotation was conducted through a dedicated team of six members, divided into two groups, and to ensure reliability and consistency, the same set of data was annotated by each of these teams. Further, to understand the agreement level mentioned in 3.2, we have used Cohen’s Kappa ($\kappa$) to measure inter-annotator agreement, achieving strong consensus scores (0.91 for binary and 0.83 for multi-class classification). Discussions were conducted for the disagreements to maintain the high-quality labelling. The annotators were briefed on the use of hate speech and derogatory language present in the comments. The workload for the annotation work was managed through a brief schedule.
Dataset Limitations and Representation:
We intended to have pan-India coverage, but the dataset reflects the national discourse on the social media platforms. As noted in Section 3.3, regions such as the North East (e.g., Tripura, Meghalaya) and Central Indian states are underrepresented compared to North-India or South-India. This imbalance is attributed to media marginalization and possibly geographic or digital barriers.
Intended Use and Content Warning:
The IndRegBias dataset contains examples of hate speech, biases, and offensive language targeting specific Indian communities. This content is solely for advanced research in bias detection and safety alignment for Large Language Models. We strictly condemn the views expressed in these comments and advise researchers to exercise caution and implement appropriate content warnings when working with or displaying this data.
Acknowledgments
We would like to express our gratitude to the annotators for their dedicated work in building the dataset: Sarabeshwar Balaji, Dishank K, Subrat Kumar Pattanayak, Akash Dutta, and Shaswat.